# Exploring the role of accretion disk geometry in shaping black hole shadows

### v1

@@ 4.1 Optically thin disks
Note that for $\psi_{0}>40^{\circ}$, the observer-black hole axis penetrates the disk, so there is no strictly dark region with zero flux inside the critical impact parameter. Nevertheless, a relatively dim area remains visible, which still corresponds to the black hole shadow in a generalized sense. It is also worth noting that the maximum observed flux decreases as $\psi_{0}$ increases. This behavior results from our assumption of a fixed mass accretion rate $\dot{M}$ across all disk models. With increasing $\psi_{0}$, the disk becomes geometrically thicker, and the emitted radiation is distributed over a larger volume. As a consequence, the energy output per unit volume-and thus the peak intensity detected by the observer-becomes smaller.
@@
The observed flux along other directions, i.e., at different values of $\alpha$, exhibits behavior intermediate between those discussed along the $z^{\prime}$ and $y^{\prime}$ axes, and the underlying analysis remains qualitatively similar. In Fig. 11, we present the observed images for optically thin disks with an inclination angle of $\theta_{0}=50^{\circ}$, illustrating the dependence on both the inner radius $r_{\rm in}$ and the half-opening angle $\psi_{0}$. In contrast, Fig. 12 shows the corresponding results for $\theta_{0}=20^{\circ}$. Note that the disk flows in the two cases exhibit different dynamics: in the former, the motion is characterized by moderate infall with $\kappa_{\rm ff}=0.5$ and low angular velocity $\kappa_{\rm K}=0.1$, while in the latter, the flow moves nearly with Keplerian angular velocity, corresponding to $\kappa_{\rm ff}=0.1$ and $\kappa_{\rm K}=0.9$. Note that for $\psi_{0}=60^{\circ}$ in Fig. 11, one can observe cross-shaped lines separating regions of differing brightness. These lines correspond to the angular boundaries of the disk-the edges of the emitting region in polar angle-which become visible when the observer-black hole axis is covered by the disk.
@@ 4.2 Optically thick disks
For a line of sight along a fixed $\alpha$ direction, whether the photon ring (peak) appears can be desided on on the condition derived from the analytical expression (15). Specifically, if the segment with the minimum inclination angle along this line of sight exceeds the right-hand side of Eq. (15), the photon ring becomes visible. Conversely, if the photon ring is absent, the lensing ring typically does not appear either, since it occurs at larger apparent radii. Another convenient criterion for identifying the presence of the photon ring (and the lensing ring) is to examine the inner edge of the direct image. If the inner edge lies at an impact parameter smaller than the critical value $b_{c}=3\sqrt{3}M$, then neither the photon ring nor the lensing ring will appear in the observed image.
@@ 4.3 Partially optically thick disks
Figure 16: Observed images for partially optically thick disks with an inclination angle $\theta_{0}=20^{\circ}$, an inner radius of $r_{\rm in}=2.3M$, a half-opening angle $\psi_{0}=10^{\circ}$ and varying effective absorption coefficients. All disks assume $\kappa_{\rm ff}=0.5$ and $\kappa_{\rm K}=0.3$. Two limiting cases-the optically thin and optically thick regimes-are also presented for comparison. The flux distributions are normalized by their respective maximum values.
@@ 5 Conclusions and Discussion
A central component of our approach was the use of transfer functions associated with these segments. These transfer functions play two critical roles: they determine the inner boundaries of the direct image, lensing ring, and photon ring; and they govern the width of the lensing ring in the case of geometrically thick disks. Both aspects are essential for understanding the structure of black hole shadows. One crucial outcome of this analysis is that, as the half-opening angle $\psi_{0}$ increases, the lensing ring systematically broadens and gradually approaches the photon ring. This behavior is a generic feature of optically thin disks. However, the initial width of the lensing ring primarily depends on the disk inclination angle $\theta_{0}$ and the inner radius $r_{\rm in}$, while the rate at which it broadens with increasing $\psi_{0}$ is influenced by the observer’s viewing direction $\alpha$. This dependence arises because a geometrically thicker disk includes emitting segments spanning a broader range of inclination angles $\varphi_{0}$. Specifically, segments with smaller $\varphi_{0}$ tend to produce lensing rings with more extended outer edges, as revealed by the transfer function in Fig. 4, whereas segments with larger $\varphi_{0}$ generate lensing rings located closer to the photon ring. Moreover, for a given disk configuration, the range of contributing $\varphi_{0}$ segments depends on the observer’s viewing direction $\alpha$, resulting in variations in the lensing ring width across different regions of the image plane.
@@
The above results provide a concrete resolution to the differing conclusions presented in Refs. Gralla:2019xty and Narayan:2019imo regarding the observability and significance of the photon ring in the images of optically thin accretion flows around a Schwarzschild black hole. Both conclusions are valid within their respective contexts: Gralla et al. Gralla:2019xty focus on geometrically thin accretion disks, while Narayan et al. Narayan:2019imo consider spherical accretion flows. Our analysis reveals that as the disk transitions from geometrically thin to spherical, the width of the photon ring remains nearly constant, typically $\Delta\lesssim 0.1M$. In contrast, the width of the lensing ring grows significantly-with variations typically on the order of $M$-and gradually bridges the photon ring and the direct image. From an observational perspective, it is therefore the lensing ring that provides a prominent and evolving signature as the disk geometry varies. This finding supports the viewpoint presented in Ref. Gralla:2019xty , namely that the lensing ring can provide a more prominent observational feature than the photon ring. Nevertheless, this does not diminish the physical importance of the photon ring. On the contrary, the photon ring precisely determines the innermost edge of the lensing ring and effectively defines the outer boundary of the black hole shadow, particularly in the case of geometrically thick accretion disks. Its presence is foundational to the overall image structure, even though its flux contribution may be subdominant or observationally blended with that of the lensing ring.
@@
In the optically thick regime, the observed flux is predominantly determined by the direct image, which corresponds to the first intersection between the light ray and the emitting surface when tracing backward from the observer. Whether the lensing ring and photon ring can appear is dictated by the observer’s viewing direction and the inclination angle of the emitting segment, as quantified by the analytical criterion given in Eq. (15). A practical criterion is that if the inner edge of the direct image-obtainable from the transfer function-lies beyond the critical impact parameter $b_{c}=3\sqrt{3}$, then neither the photon ring nor the lensing ring will appear in the observed image. Using this criterion, one can infer the qualitative structure of optically thick images from their optically thin counterparts.
@@
At last, we note that although this work focused on the Schwarzschild black hole, the method can be readily generalized to other spherically symmetric spacetimes, such as those arising in alternative theories of gravity or semiclassical quantum gravity models Capozziello:2011et ; Battista2023 ; Casadio:2025lkn ; Wang:2025fmz ; Li:2025zdt ; Tan:2024jkj . This generalization is particularly relevant given that many modified gravity scenarios predict deviations from the Schwarzschild geometry while maintaining spherical symmetry Vagnozzi:2022moj ; Wang:2022ews ; Cao:2024kht ; Wang:2023rjl ; Li:2024abk ; Gogoi:2025rcn . Such extensions may provide a useful framework for exploring potential observational signatures that distinguish general relativity from its alternatives.
@@ Appendix B Radiation Flux Formulas
The total viscous torque exerted by material at radii $>r$ on the material at radius $r$ is obtained by integrating the stress over the vertical extent of the disk:
@@
The first term in the square brackets represents the radial transport of rotational energy and is determined by boundary conditions upon integration. The second term, $-G\frac{{\rm d}\Omega}{{\rm d}r}\,{\rm d}r$, represents the local rate of mechanical energy dissipation. Assuming the dissipated energy $Q^{+}$ is radiated from both the upper and lower surfaces of the disk, the radiative flux per unit surface area is given by
@@
For slim disks or ADAF, Eq. (35) is no longer valid, since a significant fraction-or even the majority-of the dissipated energy $Q^{+}$ is transported inward by advective processes. Moreover, the orbital motion is no longer strictly Keplerian. To simplify our analysis, we adopt the following approximation for optically thick disks:
@@
we adopt Eq. (33) as an approximation for $\nu\Sigma$ in the main text. Here, we assume that viscous dissipation is the sole source of energy in the flow. For optically thin disks, the relevant quantity is the dissipated energy per unit volume:
@@
The parameter $\epsilon$ in Eqs. (36) and (38) characterizes the efficiency of radiative cooling-that is, the fraction of dissipated energy that is locally radiated. For a classical thin disk, we have $\epsilon=1$. In contrast, for ADAF, radiative cooling is highly inefficient, and $\epsilon\ll 1$. Since we are primarily interested in the envelope of the flux profile, and $\epsilon$ contributes only as an overall normalization factor, we set $\epsilon=1$ in our numerical calculations. In this situation, we have
$q^{-}=\frac{{\rm d}\Omega}{{\rm d}r}\frac{1}{4\pi\psi_{0}}\left(-\Omega+\frac{r^{2}_{\rm in}\Omega\big{|}_{r_{\rm in}}}{r^{2}}\right)\,.$ (39)
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants No. 12405063).

### v4

@@ 4.1 Optically thin disks
Note that for $\psi_{0}>40^{\circ}$, the observer-black hole axis penetrates the disk, so there is no strictly dark region with zero flux inside the critical impact parameter. Nevertheless, a relatively dim area remains visible, which still corresponds to the black hole shadow in a generalized sense. It is also worth noting that the maximum observed flux decreases as $\psi_{0}$ increases. This behavior results from our assumption of a fixed mass accretion rate $\dot{M}$ across all disk models. With increasing $\psi_{0}$, the disk becomes geometrically thicker, and the emitted radiation is distributed over a larger volume. As a consequence, the energy output per unit volume—and thus the peak intensity detected by the observer—becomes smaller.
@@
The observed flux along other directions, i.e., at different values of $\alpha$, exhibits behavior intermediate between those discussed along the $z^{\prime}$ and $y^{\prime}$ axes, and the underlying analysis remains qualitatively similar. In Fig. 11, we present the observed images for optically thin disks with an inclination angle of $\theta_{0}=50^{\circ}$, illustrating the dependence on both the inner radius $r_{\rm in}$ and the half-opening angle $\psi_{0}$. In these images, the clearly visible bright ring corresponds to the lensing ring. Whether an individual photon ring can be distinguished depends on the specific configuration. Several qualitative trends can be identified: First, for a fixed inner radius, increasing the half-opening angle causes the inner edge of the lensing ring to gradually approach the photon ring, while its outer edge progressively merges with the region of direct emission.
Second, for a fixed and relatively small half-opening angle, increasing the inner radius makes the photon ring more prominent. This occurs because the inner boundary of the lensing ring shifts outward, moving further away from the photon ring. Third, for larger half-opening angles, a relatively dim region consistently appears, enclosed by the photon ring. These features can be well understood in terms of the general behavior of the transfer functions discussed earlier.
Similar features can also be observed in Fig. 12, where we present the observed images for optically thin disks with an inclination angle of $\theta_{0}=20^{\circ}$. Note that the disk flows in the two cases (cf. Fig. 11 and 12) exhibit different dynamics: in the former, the motion is characterized by moderate infall with $\kappa_{\rm ff}=0.5$ and low angular velocity $\kappa_{\rm K}=0.1$, while in the latter, the flow moves nearly with Keplerian angular velocity, corresponding to $\kappa_{\rm ff}=0.1$ and $\kappa_{\rm K}=0.9$. Therefore, a pronounced asymmetry in the observed flux can be clearly seen in Fig. 12. Note that for $\psi_{0}=60^{\circ}$ in Fig. 11, one can observe cross-shaped lines separating regions of differing brightness. These lines correspond to the angular boundaries of the disk—the edges of the emitting region in polar angle—which become visible when the observer-black hole axis is covered by the disk.
@@ 4.2 Optically thick disks
For a line of sight along a fixed $\alpha$ direction, whether the photon ring (peak) appears can be decided on on the condition derived from the analytical expression (16). Specifically, if the segment with the minimum inclination angle along this line of sight exceeds the right-hand side of Eq. (16), the photon ring becomes visible. Conversely, if the photon ring is absent, the lensing ring typically does not appear either, since it occurs at larger apparent radii. Another convenient criterion for identifying the presence of the photon ring (and the lensing ring) is to examine the inner edge of the direct image. If the inner edge lies at an impact parameter smaller than the critical value $b_{c}=3\sqrt{3}M$, then neither the photon ring nor the lensing ring will appear in the observed image.
@@ 4.3 Partially optically thick disks
Figure 16: Observed images for partially optically thick disks with an inclination angle $\theta_{0}=20^{\circ}$, an inner radius of $r_{\rm in}=2.3M$, a half-opening angle $\psi_{0}=10^{\circ}$ and varying effective absorption coefficients. All disks assume $\kappa_{\rm ff}=0.5$ and $\kappa_{\rm K}=0.3$. Two limiting cases—the optically thin and optically thick regimes—are also presented for comparison. The flux distributions are normalized by their respective maximum values.
@@ 5 Conclusions and Discussion
A central component of our approach was the use of transfer functions associated with these segments. These transfer functions play two critical roles: they determine the inner boundaries of the direct image, lensing ring, and photon ring; and they govern the width of the lensing ring in the case of geometrically thick disks. Both aspects are essential for understanding the structure of black hole shadows. One crucial outcome of this analysis is that, as the half-opening angle $\psi_{0}$ increases, the lensing ring systematically broadens and gradually approaches the photon ring. This behavior is a generic feature of optically thin disks. However, the initial width of the lensing ring primarily depends on the disk inclination angle $\theta_{0}$ and the inner radius $r_{\rm in}$, while the rate at which it broadens with increasing $\psi_{0}$ varies with the image-plane polar angle $\alpha$. This dependence arises because a geometrically thicker disk includes emitting segments spanning a broader range of inclination angles $\varphi_{0}$. Specifically, segments with smaller $\varphi_{0}$ tend to produce lensing rings with more extended outer edges, as revealed by the transfer function in Fig. 4, whereas segments with larger $\varphi_{0}$ generate lensing rings located closer to the photon ring. Moreover, for a given disk configuration, the range of contributing $\varphi_{0}$ segments varies with the image-plane polar angle $\alpha$, resulting in variations in the lensing ring width across different regions of the image plane.
@@
The above results provide a concrete resolution to the differing conclusions presented in Refs. Gralla:2019xty and Narayan:2019imo regarding the observability and significance of the photon ring in the images of optically thin accretion flows around a Schwarzschild black hole. Both conclusions are valid within their respective contexts: Gralla et al. Gralla:2019xty focus on geometrically thin accretion disks, whereas Narayan et al. Narayan:2019imo examine spherical accretion flows. In fact, our study bridges the gap between these two scenarios. Our analysis demonstrates that as the disk transitions from a geometrically thin configuration to a spherical flow, the width of the photon ring remains nearly constant, typically $\Delta\lesssim 0.1M$. In contrast, the width of the lensing ring grows significantly—with variations typically on the order of MM—and gradually bridges the photon ring and the direct image. From an observational perspective, it is therefore the lensing ring that provides a prominent and evolving signature as the disk geometry varies. This finding supports the viewpoint presented in Ref. Gralla:2019xty , namely that the lensing ring can provide a more prominent observational feature than the photon ring. Nevertheless, this does not diminish the physical importance of the photon ring. On the contrary, the photon ring precisely determines the innermost edge of the lensing ring and effectively defines the outer boundary of the black hole shadow, particularly in the case of geometrically thick accretion disks. Its presence is foundational to the overall image structure, even though its flux contribution may be subdominant or observationally blended with that of the lensing ring.
@@
In the optically thick regime, the observed flux is predominantly determined by the direct image, which corresponds to the first intersection between the light ray and the emitting surface when tracing backward from the observer. Whether the lensing ring and photon ring can appear is determined by the image-plane polar angle $\alpha$ and the inclination angle of the emitting segment, as quantified by the analytical criterion given in Eq. (16). A practical criterion is that if the inner edge of the direct image—obtainable from the transfer function—lies beyond the critical impact parameter $b_{c}=3\sqrt{3}$, then neither the photon ring nor the lensing ring will appear in the observed image. Using this criterion, one can infer the qualitative structure of optically thick images from their optically thin counterparts.
@@
At last, we note that although this work focused on the Schwarzschild black hole, the method can be readily generalized to other spherically symmetric spacetimes, such as those arising in alternative theories of gravity or semiclassical quantum gravity models Capozziello:2011et ; Battista2023 ; Casadio:2025lkn ; Wang:2025fmz ; Li:2025zdt ; Tan:2024jkj . This generalization is particularly relevant given that many modified gravity scenarios predict deviations from the Schwarzschild geometry while maintaining spherical symmetry Vagnozzi:2022moj ; Wang:2022ews ; Cao:2024kht ; Wang:2023rjl ; Li:2024abk ; Gogoi:2025rcn ; Lim:2025cne . Such extensions may provide a useful framework for exploring potential observational signatures that distinguish general relativity from its alternatives.
@@ Appendix B Radiation Flux Formulas
The total viscous torque exerted by material at radii >r>r on the material at radius $r$ is obtained by integrating the stress over the vertical extent of the disk:
@@
The first term in the square brackets represents the radial transport of rotational energy and is determined by boundary conditions upon integration. The second term, $G\frac{{\rm d}\Omega}{{\rm d}r}\,{\rm d}r$, represents the local rate of mechanical energy dissipation. Assuming the dissipated energy $Q^{+}$ is radiated from both the upper and lower surfaces of the disk, the radiative flux per unit surface area is given by
@@
For slim disks or ADAF, Eq. (36) is no longer valid, since a significant fraction—or even the majority—of the dissipated energy $Q^{+}$ is transported inward by advective processes. Moreover, the orbital motion is no longer strictly Keplerian. To simplify our analysis, we adopt the following approximation for optically thick disks:
@@
and we adopt Eq. (34) as an approximation for $\nu\Sigma$ in the main text. Here, we assume that viscous dissipation is the sole source of energy in the flow.
For optically thin disks, the relevant quantity is the dissipated energy per unit volume:
@@
The parameter $\epsilon$ in Eqs. (37) and (39) characterizes the efficiency of radiative cooling—that is, the fraction of dissipated energy that is locally radiated. For a classical thin disk, we have $\epsilon=1$. In contrast, for ADAF, radiative cooling is highly inefficient, and $\epsilon\ll 1$. Since we are primarily interested in the envelope of the flux profile, and $\epsilon$ contributes only as an overall normalization factor, we set $\epsilon=1$ in our numerical calculations. In this situation, we have
$q^{-}=\frac{{\rm d}\Omega}{{\rm d}r}\frac{\dot{M}}{4\pi\psi_{0}}\left(-\Omega+\frac{r^{2}_{\rm in}\Omega\big|_{r_{\rm in}}}{r^{2}}\right)\,.$ (40)
@@ Acknowledgements.
We thank the anonymous referee for insightful comments and valuable suggestions, which have helped improve the clarity and quality of this manuscript. This work is supported by the National Natural Science Foundation of China (Grants No. 12405063).